OpenStudy (loser66):

Find all e. values and e. vectors of A = \(\left[\begin{matrix}0&1\\0&0\end{matrix}\right]\) I got \(\lambda =0\) double root and e. vectors are \(\left(\begin{matrix}1\\0\end{matrix}\right)\) and \(left(\begin{matrix}1\\1\end{matrix}\right)\) Please, tell me whether they are correct?

OpenStudy (loser66):

@dumbcow

OpenStudy (dumbcow):

the 2nd e vector is (1,1) ?

OpenStudy (dumbcow):

i get \[\lambda = 0\] e vectors \[\left(\begin{matrix}0 \\ 0\end{matrix}\right) , \left(\begin{matrix}1 \\ 0\end{matrix}\right)\]

OpenStudy (loser66):

how to get (0,0) ? @dumbcow

OpenStudy (dumbcow):

well \[A \left(\begin{matrix}0 \\ 0\end{matrix}\right) = 0 \left(\begin{matrix}0 \\ 0\end{matrix}\right) \] that makes it an eigenvector right?

OpenStudy (loser66):

yes

OpenStudy (loser66):

I can't guess, have to have logic on it can I take A * (1,0) =(0,0) this + (1,0) is basis of eigen space?

OpenStudy (dumbcow):

not sure, i haven't done this in awhile i forgot what to do with eigen spaces?

OpenStudy (loser66):

when I go on the normal way, it's trivial that (1,0) is a eigenvector, but it is a 2x2 matrix with double root on eigenvalue, I think we still have another eigenvector.

OpenStudy (loser66):

Ok, forget it. I think I overtake the problem. !! Thank you.